The potential difference between the plates of a parallel plate capacitor is 45 V and the electric field between the plates has a strength of 850 V/m. If the plate area is 4.0
Answer #1

use:
V=E*d
45 = (850)*d
d=0.053 m

Area, A = 4*10^-2 m^2

C= ebsolene0 * A/d
=(8.854*10^-12) * (4*10^-2) / (0.053)
= 6.7*10^-12 F

Ans: B
Answer #2

(B) 6.7
Answer #3

A) 6.7
Answer #4

(B) 6.7
Answer #5

B) 6.7
Answer #6

A) 6.7